OpenStudy (anonymous):

how do i solve 340/1 times 41/68 times 2/25

OpenStudy (anonymous):

do you want the work or justa nswer

OpenStudy (anonymous):

both

OpenStudy (anonymous):

162/1 = 162

OpenStudy (anonymous):

29/54 and 2/9 have a common denominator of 54 so just multiply the 2 by 6 and the 9 by 6 and you get 18/54

OpenStudy (anonymous):

it comes out to be 58/3

OpenStudy (anonymous):

ans is 18

OpenStudy (anonymous):

you just keep finding common denominators

OpenStudy (anonymous):

@Akpellet_math no you are wrong

OpenStudy (anonymous):

ohhhhh,,,yeah i m sorry it is 19.3.....

OpenStudy (anonymous):

which is equal to my answer

OpenStudy (anonymous):

mistakingly i took 27/54 istead of 29/54

OpenStudy (anonymous):

hmmm,..@ilikephysics2 u r correct

OpenStudy (anonymous):

you are too

OpenStudy (anonymous):

hey did u get the ans. of my ques>?

OpenStudy (anonymous):

answer

OpenStudy (anonymous):

82/5

OpenStudy (anonymous):

the answer is 16.4

OpenStudy (anonymous):

yes equivalent to mine

OpenStudy (anonymous):

very simple way to solve it : 340*41/68=205 205*2/25=16.4 i dont know, guys, why you are using such complicated ways